OpenStudy (anonymous):

Find the sum of the first 9 terms in the arithmetic sequence 45, 40, 35, 30, …

OpenStudy (ajprincess):

what is the first term and common difference here?

OpenStudy (anonymous):

45 -5

OpenStudy (ajprincess):

ya good:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

...

OpenStudy (ajprincess):

the formula to find the nth term is a_n=a_1+(n-1)d sousing this can you find the 9th term

OpenStudy (ajprincess):

otherwise directly use this formula to find the sum. S_n=(2a_1+(n-1)d)*(n/2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

9=45+(9-1)-5

OpenStudy (ajprincess):

\(a_9=45+(9-1)*(-5)\) \(=?\)

OpenStudy (anonymous):

48

OpenStudy (ajprincess):

what is 9-1=?

OpenStudy (anonymous):

8

OpenStudy (ajprincess):

ya:) 8*(-5)=?

OpenStudy (anonymous):

-48

OpenStudy (ajprincess):

nope:(

OpenStudy (anonymous):

-40

OpenStudy (ajprincess):

ya:) -40+45=?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Thank you the answer is 5

OpenStudy (ajprincess):

ya:) that is ur 9th term. nw u knw the first term and last term of the ap whose sum u need to find

OpenStudy (ajprincess):

the sum=(n/2)*(a_1+a_9)

OpenStudy (ajprincess):

did u find the sum?

OpenStudy (anonymous):

I dont think I need the sum

OpenStudy (anonymous):

oh nvm I do need it

OpenStudy (anonymous):

45+40+35+30+25+20+15+10+5

OpenStudy (anonymous):

225

OpenStudy (ajprincess):

the question is to find the sum. to do that u can use the formula. S_n=(n/2)*(a_1+a_n) =(9/2)(45+5) =(9/2)(50) =9*25 =225 u dnt have to write all the terms nd add them ya u r right:)